# The Language of Approval: Identifying the Drivers of Positive Feedback Online

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; Under Review; 2025ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1.1. Opportunities and Challenges in Positive Feedback Mechanisms
The power of these reward mechanisms extends far beyond simple content ranking. Prior research demonstrates that positive feedback creates reinforcement loops that actively shape future behavior. Being highly upvoted or receiving awards on Reddit encourages users to post content that receives even more positive feedback than they would have otherwise (Lambert et al., 2025). Similarly, creator recognition features such as “creator hearts” on YouTube improve engagement and comment visibility (Choi et al., 2025). This pattern reflects broader principles of positive reinforcement validated across offline contexts, from workplace productivity to child development (Gray, 2010; Wei and Yazdanifard, 2014; Athalye et al., 2018). Reward systems do not merely reflect community preferences—they actively incentivize behaviors that align with perceived community values (Weld et al., 2021; Goyal et al., 2026).
@@
However, a fundamental gap limits progress towards building such systems: we lack a causal understanding of what drives positive feedback. Current research falls into two categories, each with its own limitations. First, surveys of moderators and users identify what communities say they value (Lambert et al., 2024; Weld et al., 2022, 2024), but surveys may reflect aspirations rather than realized behavioral patterns in practice. Second, descriptive and correlational studies identify attributes in content that are associated with higher engagement (Bär et al., 2023; Papakyriakopoulos et al., 2023; Voinea et al., 2024; Goyal et al., 2026), but correlation conflates language with author reputation, timing, and visibility. We note that this distinction is crucial: knowing that popular posts tend to use specific linguistic pattern is a reflective observation, providing limited guidance for actionable proactive interventions. Without isolating causal effects, we cannot build reliable detection or guidance tools, or explain why certain interventions succeed or fail.
@@ 1.3.1. Methods:
To answer these research questions, we conduct a large scale empirical study using 11 million comments from 100 communities on Reddit (or subreddits) to make three key contributions in this work. First, using a quasi-experimental causal inference framework, we identify the attributes from a rich set of linguistic, topical, and semantic features that influence the likelihood of a post receiving positive feedback. Second, we use predictive modeling on our feature set to model the likelihood of a post being rewarded by the community. Finally, we conduct an audit of existing community guidelines to identify what kind of rules are currently prescribed, and whether they include rules to teach users how to make desirable contributions in context of our findings.
@@ 2.1. Proactive Community Governance
This imbalance exists even though both theoretical foundations and empirical evidence supporting the effectiveness of positive reinforcement mechanisms. The psychological foundations for positive reinforcement date back to early behavioral research by Ferster and Skinner (1957), which established that providing positive stimuli following desired behaviors effectively encourages their repetition. This principle has been validated across diverse offline contexts, from workplace productivity (Perryer et al., 2016) to educational settings (Steinberg et al., 1992) and sports performance (Wiese et al., 1991). Importantly, research demonstrates that positive reinforcement can operate both directly on individuals and vicariously through observational learning (Bandura and Barab, 1971; Kazdin, 1973), suggesting that highlighting exemplary behavior can influence broader community norms.
In online contexts, HCI researchers have increasingly recognized the potential of positive reinforcement for improving community dynamics. Kraut et al. (2011) identify positive feedback as a key mechanism for encouraging sustained user participation, while Kiesler et al. (2012) emphasize how highlighting norm-adhering behavior can vicariously reinforce similar patterns among community members. This theoretical foundation has gained empirical support through several platform-specific studies. For example, Wang and Diakopoulos (2021) demonstrated that introducing a simple badge system for New York Times commenters significantly improved both comment quality and frequency, with particularly strong effects for newcomers. Gurjar et al. (2022) showed that when content posted by social media users receives sudden viral attention, the users significantly increase posting frequency and alter their content to more closely match the viral post. Lambert et al. (2025) showed that positive feedback mechanisms on Reddit causally improve users’ future content quality and reduce removal rates, while Choi et al. (2025) showed that using “creator hearts” as a positive reinforcement strategy for desirable comments on YouTube improves engagement and comment visibility.
@@ 2.2. Content Quality Assessment and Modeling Community Feedback Mechanisms
On the other hand, many studies focus on empirical approaches to predicting content success, operationalize certain measures, or explain community dynamics. For example, researchers have developed models using temporal patterns (Szabo and Huberman, 2010) and engagement cascades (Cheng et al., 2014). Mitra and Gilbert (2014) studied the factors which led to successfully funding a crowdfunding projects on Kickstarter, but do not offer causal claims. Lambert et al. (2022) studied what attributes in content enable conversations on Reddit to be resilient to adverse events such as content removals. Similarly, the framework and models introduced by  Bao et al. (2021) have been used widely for measuring desirable contributions, although it captures only a subset of factors that drive positive feedback, i.e., prosociality. The work of Atcheson et al. (2024) focuses on language at an averaged sentence embedding level to demonstrate that online communities face a trade-off between growth and linguistic distinctiveness. This further suggests that community evolution fundamentally reshapes the linguistic landscape that can determine what content receives positive reception. These approaches primarily focus on prediction and measuring correlation rather than understanding causal mechanisms.
Prior research that has focused on answering causal questions have studied downstream effects of receiving positive/negative feedback (Chandrasekharan et al., 2022; Papakyriakopoulos et al., 2023) or high engagement (Gurjar et al., 2022), which means an in-depth understanding of what actually causes this feedback is missing. Similarly, research on mental health outcomes has used quasi-experimental designs to examine the effectiveness of the Papageno effect (Domaradzki, 2021) for suicidal ideation on Twitter (now X) (Yuan et al., 2023). Different from our work, these studies focus on behavioral outcomes rather than linguistic mechanisms. More similar to our setting, Zhang et al. (2018) applied causal inference to study conversational failure, i.e., isolate the effects of early warning signals on conversations that end in undesirable outcomes.
@@ 3. Data Curation, Methods, and Study Design
Our goal is to quantify how observable post-level attributes—e.g., when the submission is made, who made it (i.e., the author’s baseline activity, tenure in the community, and prior karma), and its content—shape the likelihood of receiving positive feedback. Formally, we ask: By how much does the probability of earning a high score, an award, or a gold change when a given attribute increases or decreases, holding all other observed factors constant? Because a single post cannot exist simultaneously with and without that attribute, the counterfactual outcome is unobserved. We therefore adopt a selection-on-observables (Cochran, 1968; Eggers and Hainmueller, 2009; Sekhon, 2009; Stuart, 2010) framework for causal inference. In our setting, this means that, conditional on observed author history/activity and on community/timing fixed effects (i.e., the observed covariates), we assume that there are no unmeasured factors that jointly affect the focal linguistic attribute of interest (i.e., the treatment) and the likelihood of positive feedback (i.e., the outcome)—so the remaining variation in that attribute is as good as random with respect to the outcomes.
Our data comes from the Pushshift archives of Reddit data (Baumgartner et al., 2020) from May 1, 2020 to Sep 30, 2020, mirroring the study window of Lambert et al. (2025). We chose this window to maintain consistency with prior work on studying positive reinforcement on Reddit, and to ensure that all commenting and voting activity in these subreddits has stabilized by the time of this study. This ensures that the community feedback we model is representative. Since our research did not involve any user participation, all data was publicly available on Reddit, and we ensured to maintain anonymity in our data to ensure user privacy, our work did not require approval from an institutional review board (IRB). We describe our feedback outcomes and candidate pool construction in the sections below.
@@ 3.1. Positive Feedback Outcomes
To complement this broad signal, we also consider two paid forms of positive feedback: awards and gold. These signals are comparatively sparse in practice but carry a higher bar of endorsement because they incur a monetary cost. Reddit Gold was a popular yet infrequently used mechanism (Wang and Diakopoulos, 2021), and was removed in 2023 (Weatherbed, 2024). Awards were also removed in 2023111https://reddit.com/r/reddit/comments/14ytp7s/reworking_awarding_changes_to
_awards_coins_and/
@@
Note that, consistent with prior work (Lambert et al., 2025; Goyal et al., 2026), we interpret score, awards, and gold as a reaction signal of approval by the community rather than a guarantee of accuracy or quality. There is a possibility that these mechanisms do not necessarily represent a high-quality contribution, and that there are more important signals of “desirability” than getting highly upvoted.
@@ 3.2. Candidate-pool Construction
Following Lambert et al. (2025), we partition our overall observation window into a baseline window and a sampling window. The overall observation window is May 1–Sep 30, 2020. The 14-day baseline spans May 1–May 14 and is used to compute pre-feedback author covariates (e.g., posting and removal rates, typical score/awards, account age). The sampling window spans May 15–Sep 30, 2020 during which candidates are constructed and outcomes are evaluated.
@@ 3.4.1. Linguistic Feature Set ($\mathbf{A}$):
  • Latent topics: We estimate topic probabilities with Latent Dirichlet Allocation (LDA) (Blei et al., 2003) using K=10K{=}10 topics (subjectively determined through coherence inspection); topic probabilities lie in $[0,1]$ and sum to 1 per post, capturing coarse content structure. LDA has been used extensively in prior social media research as it produces stable and interpretable topics (Chancellor et al., 2016; Ernala et al., 2017; Saha et al., 2024).
@@ 3.5. Feature Engineering and Preprocessing
After collecting the baseline covariates ($\mathbf{Z}$), the linguistic feature set ($\mathbf{A}$) and newcomer indicators (NEW), we perform some feature engineering and data preprocessing steps which we describe below.
@@ 3.5.1. LIWC Residual Umbrellas:
LIWC categories are hierarchical, where each coarse “umbrella” constructs such as “cogproc,” “affect,” etc. contain finer-grained children, for e.g., insight, cause for “cogproc” and posemo, negemo for “affect.” Our initial experimentation revealed that when umbrella and children terms are used together as predictors, the umbrella term is often highly correlated with the sum of its children, which leads to multicollinearity and ambiguous coefficients (e.g., unstable magnitudes or sign flips as umbrella and its child attributes compete to explain the same variance).
Therefore, to maintain interpretability, we want the model to include an umbrella-level signal that provides signal beyond what the children already capture. In order to do this, for each umbrella $U$ with children $C_{U}$ we regress $U$ on its children across posts and replace the umbrella with the residual.
Concretely, for each post $i=1,\dots,n$ we fit a linear model of the umbrella on its children:
@@ 3.6. Risk Stratification and Balance
using an AdaBoost classifier (Freund et al., 1999) to flexibly capture nonlinearities and interactions in $\mathbf{Z}$. We intentionally exclude linguistic features $\mathbf{A}$ from the risk model so that linguistic variation remains to explain residual differences within well-balanced strata. Using a treatment-style propensity score similar to a potential-outcomes framework (Rubin, 2005) would require a single well-defined treatment. We therefore do not use propensity score as our linguistic attributes are not binary “treatments” in the traditional sense of a potential outcomes framework.
We bin our outcome risk scores $r(\mathbf{Z},s)$ into deciles and retain strata with adequate overlap (minimum ten positive-feedback and control candidates). To assess covariate balance between positive and control candidates within each stratum, we compute standardized mean differences (SMD) for all covariates in $\mathbf{Z}$ and, following prior work, discard strata whose mean SMD across $\mathbf{Z}$ exceeds $0.30$ (Kampenes et al., 2007; Kiciman et al., 2018; Lambert et al., 2025) (see Figure 3). This design yields many-to-many, within-subreddit comparisons among posts with similar baseline risk of receiving positive feedback.
[…]
@@ 4.1. RQ1(a): Which attributes raise or lower the likelihood of positive feedback?
PC Odds Ratio $q$-sig Interpretation from Exemplars Implication for High Score Reward PC1 1.43 *** Narrow help–requests vs. broad engagement prompts ↑↑ Broad, discussion-generating posts PC2 0.99 - Text-only discussion vs. visual/creative showcases ↓↓ Text-only content less favored PC3 1.00 - Personal‐struggle narratives vs. creative achievements PC4 0.98 - Dense, list-like technical exposition ↓↓ Complex enumerations PC5 0.98 - Off-topic digressions / meta commentary ↓↓ Meta or digressive content PC6 0.94 *** Casual social updates vs. substantive discussion ↓↓ Brief ‘‘status’’ posts PC7 1.03 - Enthusiastic, community-oriented calls ↑↑ Collective enthusiasm PC8 0.93 *** Event-dated / time-stamped content ↓↓ Heavily time-bound posts PC9 0.97 - Literal descriptions vs. reflective commentary ↓↓ Purely literal recounting PC10 0.95 * Serendipitous “found this” shares ↓↓ Chance discoveries over curated insight

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790476isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1.1. Opportunities and Challenges in Positive Feedback Mechanisms
The power of these reward mechanisms extends far beyond simple content ranking. Prior research demonstrates that positive feedback creates reinforcement loops that actively shape future behavior. Being highly upvoted or receiving awards on Reddit encourages users to post content that receives even more positive feedback than they would have otherwise (Lambert et al., 2025b). Similarly, creator recognition features such as “creator hearts” on YouTube improve engagement and comment visibility (Choi et al., 2025). This pattern reflects broader principles of positive reinforcement validated across offline contexts, from workplace productivity to child development (Gray, 2010; Wei and Yazdanifard, 2014; Athalye et al., 2018). Reward systems do not merely reflect community preferences—they actively incentivize behaviors that align with perceived community values (Weld et al., 2021; Goyal et al., 2026a).
@@
However, a fundamental gap limits progress towards building such systems: we lack a causal understanding of what drives positive feedback. Current research falls into two categories, each with its own limitations. First, surveys of moderators and users identify what communities say they value (Lambert et al., 2024; Weld et al., 2022, 2024), but surveys may reflect aspirations rather than realized behavioral patterns in practice. Second, descriptive and correlational studies identify attributes in content that are associated with higher engagement (Bär et al., 2023; Papakyriakopoulos et al., 2023; Voinea et al., 2024; Goyal et al., 2026a), but correlation conflates language with author reputation, timing, and visibility. We note that this distinction is crucial: knowing that popular posts tend to use specific linguistic pattern is a reflective observation, providing limited guidance for actionable proactive interventions. Without isolating causal effects, we cannot build reliable detection or guidance tools, or explain why certain interventions succeed or fail.
@@ 1.3.1. Methods:
To answer these research questions, we conduct a large scale empirical study using 11 million posts from 100 communities on Reddit (or subreddits) to make three key contributions in this work. First, using a quasi-experimental causal inference framework, we identify the attributes from a rich set of linguistic, topical, and semantic features that influence the likelihood of a post receiving positive feedback. Second, we use predictive modeling on our feature set to model the likelihood of a post being rewarded by the community. Finally, we conduct an audit of existing community guidelines to identify what kind of rules are currently prescribed, and whether they include rules to teach users how to make desirable contributions in context of our findings.
@@ 2.1. Proactive Community Governance
This imbalance exists even though both theoretical foundations and empirical evidence supporting the effectiveness of positive reinforcement mechanisms. The psychological foundations for positive reinforcement date back to early behavioral research by Ferster and Skinner (1957), which established that providing positive stimuli following desired behaviors effectively encourages their repetition. This principle has been validated across diverse offline contexts, from workplace productivity (Perryer et al., 2016) to educational settings (Steinberg et al., 1992) and sports performance (Wiese et al., 1991). Importantly, research demonstrates that positive reinforcement can operate both directly on individuals and vicariously through observational learning (Bandura and Barab, 1971; Kazdin, 1973), suggesting that highlighting exemplary behavior can potentially influence broader community norms.
In online contexts, HCI researchers have increasingly recognized the potential of positive reinforcement for improving community dynamics. Kraut et al. (2011) identify positive feedback as a key mechanism for encouraging sustained user participation, while Kiesler et al. (2012) emphasize how highlighting norm-adhering behavior can vicariously reinforce similar patterns among community members. This theoretical foundation has gained empirical support through several platform-specific studies. For example, Wang and Diakopoulos (2021) demonstrated that introducing a simple badge system for New York Times commenters significantly improved both comment quality and frequency, with particularly strong effects for newcomers. Gurjar et al. (2022) showed that when content posted by social media users receives sudden viral attention, the users significantly increase posting frequency and alter their content to more closely match the viral post. Lambert et al. (2025b) showed that positive feedback mechanisms on Reddit causally improve users’ future content quality and reduce removal rates, while Choi et al. (2025) showed that using “creator hearts” as a positive reinforcement strategy for desirable comments on YouTube improves engagement and comment visibility.
@@ 2.2. Content Quality Assessment and Modeling Community Feedback Mechanisms
On the other hand, many studies focus on empirical approaches to predicting content success, operationalize certain measures, or explain community dynamics. For example, researchers have developed models using temporal and environmental patterns (Szabo and Huberman, 2010; Shimgekar et al., 2025) and engagement cascades (Cheng et al., 2014). Mitra and Gilbert (2014) studied the factors which led to successfully funding a crowdfunding projects on Kickstarter, but do not offer causal claims. Lambert et al. (2022) studied what attributes in content enable conversations on Reddit to be resilient to adverse events such as content removals. Similarly, the framework and models introduced by  Bao et al. (2021) have been used widely for measuring desirable contributions, although it captures only a subset of factors that drive positive feedback, i.e., prosociality. The work of Atcheson et al. (2024) focuses on language at an averaged sentence embedding level to demonstrate that online communities face a trade-off between growth and linguistic distinctiveness. This further suggests that community evolution fundamentally reshapes the linguistic landscape that can determine what content receives positive reception. These approaches primarily focus on prediction and measuring correlation rather than understanding causal mechanisms.
Prior research that has focused on answering causal questions have studied downstream effects of receiving positive/negative feedback (Chandrasekharan et al., 2022; Papakyriakopoulos et al., 2023) or high engagement (Gurjar et al., 2022), which means an in-depth understanding of what actually causes this feedback is missing. Similarly, research on mental health outcomes has used quasi-experimental designs to examine the effectiveness of the Papageno effect (Domaradzki, 2021) for suicidal ideation on Twitter (now X) (Yuan et al., 2023), and the effect of news exposure on well-being outcomes (Pal et al., 2026). Different from our work, these studies focus on behavioral outcomes rather than linguistic mechanisms. More similar to our setting, Zhang et al. (2018) applied causal inference to study conversational failure, i.e., isolate the effects of early warning signals on conversations that end in undesirable outcomes.
@@ 3. Data Curation, Methods, and Study Design
Our goal is to quantify how observable post-level attributes—e.g., when the submission is made, who made it (i.e., the author’s baseline activity, tenure in the community, and prior karma), and its content—shape the likelihood of receiving positive feedback. Formally, we ask: By how much does the probability of earning a high score, an award, or a gold change when a given attribute increases or decreases, holding all other observed factors constant? Because a single post cannot exist simultaneously with and without that attribute, the counterfactual outcome is unobserved. We therefore adopt a selection-on-observables framework (Cochran, 1968; Eggers and Hainmueller, 2009; Sekhon, 2009; Stuart, 2010) for causal inference. In our setting, this means that, conditional on observed author history/activity and on community/timing fixed effects (i.e., the observed covariates), we assume that there are no unmeasured factors that jointly affect the focal linguistic attribute of interest (i.e., the treatment) and the likelihood of positive feedback (i.e., the outcome)—so the remaining variation in that attribute is as good as random with respect to the outcomes.
We employ selection-on-observables because stronger causal designs are infeasible for our research. Randomized experiments (RCT) would require manipulating user content (unethical and/or ecologically invalid), while difference-in-differences (DiD) and regression discontinuity (RDiT) require specific treatment structures—discrete timing, exogenous instruments, or sharp discontinuities—that don’t exist for continuously varying linguistic choices like in our setting. Our approach requires assuming conditional independence, where after controlling for author history, timing, and community context, residual linguistic variation is as-good-as-random. While this assumption is untestable, our design represents a substantial methodological advancement over descriptive studies by isolating linguistic effects from major confounds, providing conditional causal estimates rather than correlational associations. We therefore highlight that our estimates should be interpreted as conditional causal effects under stated assumptions rather than assumption-free causal claims. See Section 7.6.3 for further discussion.
Our data comes from the Pushshift archives of Reddit data (Baumgartner et al., 2020) from May 1, 2020 to Sep 30, 2020, mirroring the study window of Lambert et al. (2025b). We chose this window to maintain consistency with prior work on studying positive reinforcement on Reddit, and to ensure that all commenting and voting activity in these subreddits has stabilized by the time of this study. This ensures that the community feedback we model is representative. Since our research did not involve any user participation, all data was publicly available on Reddit, and we ensured to maintain anonymity in our data to ensure user privacy, our work did not require approval from an institutional review board (IRB). We describe our feedback outcomes and candidate pool construction in the sections below.
@@ 3.1. Positive Feedback Outcomes
To complement this broad signal, we also consider two paid forms of positive feedback: awards and gold. These signals are comparatively sparse in practice but carry a higher bar of endorsement because they incur a monetary cost. Reddit Gold was a popular yet infrequently used mechanism (Wang and Diakopoulos, 2021), and was removed in 2023 (Weatherbed, 2024). Awards were also removed in 2023111https://reddit.com/r/reddit/comments/14ytp7s/reworking_awarding_changes_to_
awards_coins_and/
@@
Note that, consistent with prior work (Lambert et al., 2025b; Goyal et al., 2026a), we interpret score, awards, and gold as a reaction signal of approval by the community rather than a guarantee of accuracy or quality. There is a possibility that these mechanisms do not necessarily represent a high-quality contribution, and that there are more important signals of “desirability” than getting highly upvoted (see Section 7.5).
@@ 3.2. Candidate-pool Construction
Following Lambert et al. (2025b), we partition our overall observation window into a baseline window and a sampling window. The overall observation window is May 1–Sep 30, 2020. The 14-day baseline spans May 1–May 14 and is used to compute pre-feedback author covariates (e.g., posting and removal rates, typical score/awards, account age). A 14-day baseline window is standard in prior social-media based causal inference works (Yuan et al., 2023; Lambert et al., 2025b). The sampling window spans May 15–Sep 30, 2020 during which candidates are constructed and outcomes are evaluated.
@@ 3.4.1. Linguistic Feature Set ($\mathbf{A}$):
  • Latent topics: We estimate topic probabilities with Latent Dirichlet Allocation (LDA) (Blei et al., 2003) using $K\in\{5,10,15,20,25\}$ topics; topic probabilities lie in $[0,1]$ and sum to 1 per post, capturing coarse content structure. LDA has been used extensively in prior social media research as it produces stable and interpretable topics (Chancellor et al., 2016; Ernala et al., 2017; Saha et al., 2024).
@@ 3.5. Feature Engineering and Preprocessing
After collecting the baseline covariates ($\mathbf{Z}$), the linguistic feature set ($\mathbf{A}$) and the newcomer indicator (NEW), we perform some feature engineering and data preprocessing steps which we describe below.
@@ 3.5.1. LIWC Residual Umbrellas:
LIWC categories are hierarchical, where each coarse “umbrella” constructs such as “cogproc,” “affect,” etc. contain finer-grained children, for e.g., insight, cause for “cogproc” and posemo, negemo for “affect.” Our initial experimentation revealed that when umbrella and children terms are used together as predictors, the umbrella term is often highly correlated with the sum of its children, which leads to multicollinearity and ambiguous coefficients (e.g., unstable magnitudes or sign flips as umbrella and its child attributes compete to explain the same variance). Therefore, to maintain interpretability and robustness in our methodology, we want the model to include an umbrella-level signal that provides signal beyond what the children already capture.
To do so, for each umbrella $U$ with children $C_{U}$ we regress $U$ on its children across posts and replace the umbrella with the residual. Concretely, for each post $i=1,\dots,n$ we fit a linear model of the umbrella on its children:
@@ 3.6. Risk Stratification and Balance
using an AdaBoost classifier (Freund et al., 1999). Prior causal inference work has used simpler models such as the averaged perceptron learning algorithm (De Choudhury et al., 2016) and logistic regression (Yuan et al., 2023) for risk stratification and propensity score matching, however recent works such as Lambert et al. (2025b) use Adaboost as it flexibly capture nonlinearities and interactions in the covariates, and we therefore use Adaboost in our work. We intentionally exclude linguistic features $\mathbf{A}$ from the risk model so that linguistic variation remains to explain residual differences within well-balanced strata. Using a treatment-style propensity score similar to a potential-outcomes framework (Rubin, 2005) would require a single well-defined treatment. We therefore do not use propensity score as our linguistic attributes are not binary “treatments” in the sense of a potential outcomes framework.
We bin our outcome risk scores $r(\mathbf{Z},s)$ into deciles and retain strata with adequate overlap (minimum ten positive-feedback and control candidates). To assess covariate balance between positive and control candidates within each stratum, we compute standardized mean differences (SMD) for all covariates in $\mathbf{Z}$ and, following prior work, discard strata whose mean SMD across $\mathbf{Z}$ exceeds $0.30$ (Kampenes et al., 2007; Kiciman et al., 2018; Lambert et al., 2025b) (see Figure 3). This design yields many-to-many, within-subreddit comparisons among posts with similar baseline risk of receiving positive feedback.
@@ 4.1. RQ1(a): Which attributes raise or lower the likelihood of positive feedback?
PC Odds Ratio $q$-sig Interpretation from Exemplars Implication for High Score Reward
PC1 1.43 *** Narrow help–requests vs. broad engagement prompts ↑↑ Broad, discussion-generating posts
PC2 0.99 - Text-only discussion vs. visual/creative showcases ↓↓ Text-only content less favored
PC3 1.00 - Personal‐struggle narratives vs. creative achievements
PC4 0.98 - Dense, list-like technical exposition ↓↓ Complex enumerations
PC5 0.98 - Off-topic digressions / meta commentary ↓↓ Meta or digressive content
PC6 0.94 *** Casual social updates vs. substantive discussion ↓↓ Brief ‘‘status’’ posts
PC7 1.03 - Enthusiastic, community-oriented calls ↑↑ Collective enthusiasm
PC8 0.93 *** Event-dated / time-stamped content ↓↓ Heavily time-bound posts
PC9 0.97 - Literal descriptions vs. reflective commentary ↓↓ Purely literal recounting
PC10 0.95 * Serendipitous “found this” shares ↓↓ Chance discoveries over curated insight